OpenStudy (anonymous):

a small blok slides, without friction, down an inclined plane starting from rest. let sn be the distance travelled from t=(n-1) to t=(n). then (sn)/(sn+1)???

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

Let a be the constant acceleration of the motion. Let the block be released at t=0. Speed after (n-1) seconds is given by v=a(n-1) sn=vt+0.5at^2 t=1second (since sn is distance travelled from t=n-1 to t=n) v=a(n-1) , the initial speed at t=n-1 seconds So, sn=a(n-1) +a/2 =an-a/2 .....................................(1) Speed after time n seconds will be ,v= an So sn+1 = an +a/2 .....................................(2) [you may get it from the same procedure as above] sn/sn+1 = (n-0.5) / (n+0.5)

OpenStudy (anonymous):

so we can say it is equal to (2n-1)/(2n+1).... jeee thnks a bunch

OpenStudy (anonymous):

Remember n has to be greater than or equal to 1 for this result to hold good. If suppose you take n=0.1sec. Then s1 is distance travelled from t=-0.9 sec to t=0.1 sec. But motion was not present before t=0 and we get meaningless result.

OpenStudy (anonymous):

ok... thnq.... btw.... u a lecturer?? :)

OpenStudy (anonymous):

Nope, a student!

OpenStudy (anonymous):

lol.... which grade?

OpenStudy (anonymous):

currently enjoying holidays!

OpenStudy (anonymous):

I am in first year of graduation.

OpenStudy (anonymous):

what about u?

OpenStudy (anonymous):

cool....thnx a bunch dude.... me in +1... nd ya i hav anoder question... me taggin u on it.... plz do help me wid it

OpenStudy (anonymous):

no problem. Always glad to help if i can